# shapes,patterns,positioning

### v2

@@ 4.1 Double categories
We will denote $\alpha:l\to r$ for $\alpha$ a square with vertical domain $l$ and codomain $r$. If $\alpha$ is a square of $\mathbb{S}\textbf{q}(\mathcal{C})$ with horizontal domain $u$ and codomain $v$, we will denote $\alpha=[u,v]$.
@@ 4.2 Algebraic weak factorisation systems
    For any {tikzcd} square in $\mathbb{L}$ and {tikzcd} square in $\mathbb{R}$, (29)
@@
[…]
@@
Now we will state without proof Proposition 18 from Bou (23):
@@
Such lifting structures are called cofibrantly generated. It should be noted that the proof for the above proposition uses a version of Quillen’s small objects argument from Gar (08), where the axiom of choice was used. At the moment, it is not known whether a constructive proof for this argument exists.
@@ 4.3 Degenerate-preferring Kan fibrations
[…]
@@
The left vertical compatibility conditions are trivially satisfied as whenever two vertical arrows can be composed, one of them is an identity. Satisfying the left horizontal compatibility conditions agains identity squares is also trivial, thus the only squares for which it is not trivial to satisfy the left horizontal compatibility condition are those as in Equation 39.
@@
Then $s=s^{\prime}\circ s_{j}$ for some degeneracy map $s_{j}$, hence $z\circ s$ is a degenerate solution for the lifting problem against $\iota$ and therefore $\text{lift}(z\circ s\circ i,f\circ s)=z\circ s$ as $\alpha$ is degenerate-preferring. Since $z$ must be the unique lift against the identity arrow, $(\alpha,\text{lift})$ satisfies the left horizontal compatibility condition against the above squares as in Equation 39 and is a vertical morphism of ${\bf RLP}(\mathbb{D})$.
Conversely, suppose that a degenerate solution $z\circ s_{j}$ to a lifting problem as in Equation 7 exists, then the square in Equation 7 factors as the above diagram for $s=s_{j}$. Note that the right square has $z$ as unique lift. By the horizontal compatibility condition, any vertical morphism $(f,\psi)$ of ${\bf RLP}(\mathbb{D})$ therefore assigns $z\circ s_{j}$ as solution to Equation 7. Therefore any such a vertical morphism is a degenerate-preferring Kan fibration.
@@ 4.4 The double category of horn pushout sequences
In particular, in Equation 42, the sieve $S_{i}$ contains exactly two more non-degenerate simplices than $S_{i-1}$, namely, one nondegenerate $n$-simplex and one nondegenerate $n-1$-simplex. As a consequence, in Equation 41, the sieve $S_{k}$ has exactly $2k$ more non-degenerate simplices than $S_{0}$.
Now for the squares, our first attempt would be the following definition:
@@ 4.5 Functional lifts respect face squares
Note that for any $\iota\in\mathcal{L}$ with $\ell(\iota):\Lambda^{n}_{m}\hookrightarrow\Delta^{n}$, both $\Lambda^{n}_{m}$ and $\Delta^{n}$ are decidable sieves of $\Delta^{n}$ and therefore we have a horn pushout sequence of length $1$ corresponding to $[id,id]:\ell(\iota)\to\ell(\iota)$. If $(f,\psi)$ is a vertical morphism in $RLP(\mathbb{S})$, we call its lift against this sequence the lift against $\iota$.
@@
    So let $\delta:\sigma\to\tau$ be a face square and let $\tau$ have length $1$ with chosen $\iota\in\mathcal{L}$. As pullbacks of face maps are mono, the square $L(\sigma)\to\tau_{1}$ in sSet is a pullback square. By Remark 4.7, $L(\sigma)$ adds at most 2 faces, thus $\sigma$ has either length 0 or 1. If the length is 00, there is a unique lift against $\sigma$ and we are done.
@@ 4.6 Effective lifts respect degeneracy squares
In this section, we will study what happens if we respect not only face squares, but also degeneracy squares. By Lemma 4.11, we only need to consider degeneracy squares $\delta:\sigma\to\tau$ with $\tau$ of length $1$ as in Equation 48. Let $s_{j}:\Delta^{b+1}\to\Delta^{b}$ be the underlying simplicial morphism of $\delta$. It is decidable whether the index $j:\Delta^{0}\to\Delta^{b}$ factors through the inclusion $\Delta^{n}\hookrightarrow\Delta^{b}$.
@@
    If $j$ does not factor through $\Delta^{n}$, then the pullback of the diagram in Equation 48 restricts to the identity on both $\Lambda^{n}_{m}$ and $\Delta^{n}$ and, therefore, $\Lambda^{n}_{m}$ factors through $s_{j}^{*}(T_{0})$ and $\Delta^{n}$ factors through $s_{j}^{*}(T_{1})$. Therefore, for $(f,\psi)$ a vertical morphism of $RLP(\mathbb{S})$, we arrive in the same situation as in Equation 47, where the vertical domain of $\delta$ has length 1. By Lemma 4.13, both maps $s_{j}^{*}(T_{1})\to X,T_{1}\to X$ are defined by a pushout, and by the universal property of pushouts, both maps $s_{j}^{*}(T_{1})\to X$ are equal, and we respect $\delta$.
@@
      If $j^{\prime}=m$, then $\sigma$ has length 2. The first inclusion $\sigma_{1}$ is the pushout of a horn inclusion $\Lambda^{n}_{m}\hookrightarrow\Delta^{n}$ corresponding to either face $d_{j^{\prime}}$ or $d_{j^{\prime}+1}$ of $\Delta^{n+1}$, while the second inclusion $\sigma_{2}$ is the pushout of the horn inclusion $\Delta^{n+1}_{m^{*}}\hookrightarrow\Delta^{n+1}$ for $m^{*}$ the face we did not add in $\sigma_{1}$.
@@
      If $j^{\prime}\neq m$, then $\sigma$ has length 3. In this case $\sigma_{1}$ and $\sigma_{2}$ are both the pushout of a horn inclusion $\Delta^{n}_{m}\hookrightarrow\Delta^{n}$, corresponding to either face $d_{j^{\prime}}$ or $d_{j^{\prime}+1}$ of $\Delta^{n+1}$, while $\sigma_{3}$ is the pushout of the horn inclusion $\Delta^{n+1}_{m^{*}}\hookrightarrow\Delta^{n+1}$ for $m^{*}=m$ if $m<j^{\prime}$ and $m^{*}=m+1$ if $m>j^{\prime}$.
    For each $\sigma_{i}$ we need a chosen horn inclusion $\iota^{\prime}\in\mathcal{L}$ such that $\sigma_{i}$ is the pushout of $\ell(\iota^{\prime})$. The chosen horn inclusion for the final $\sigma_{i}$ (in the cases above $\sigma_{2}$ and $\sigma_{3}$ respectively) is denoted $\iota^{*}_{\delta}$, for $\iota$ as in Equation 48. Note that if we respect face squares, the lifts corresponding to the faces $d_{j^{\prime}},d_{j^{\prime}+1}$ of $\Delta^{n+1}$, will be the same as the lift against $\iota$.
@@ 5 Outlook
In Section 2, we have introduced degenerate-preferring and symmetric effective Kan fibrations, both can be given the structure of effective Kan fibrations. In BF (22), the effective Kan fibrations are constructively shown to be closed under pushforwards. This is a first step in a constructive model for HoTT based on simplicial sets. It remains an open question whether these new notions of Kan fibrations are themselves also closed under pushforwards, and can form the basis for another HoTT model.
In Section 3, we have extended the result that all simplicial algebras of an algebraic theory are Kan fibrations iff the algebraic theory is Malcev to our other notions of Kan complexes. As we mentioned in Remark 3.14, there might be an extension of this result to Malcev categories. Another direction for generalization, lies in the Malcev structure in cubical sets with connections. In Ton (92), it has been shown that cubical groups with connections are Kan. It is an open question whether this also holds for cubical Malcev algebras with connections, and whether all algebraic theories whose cubical algebras with connections are Kan are Malcev algebras.
In Section 4, we have presented proofs showing that our new notions fit into lifting algebraic weak factorisation systems. We have used Proposition 4.3, for which it is not yet known whether a constructive proof exists. We should point out that we use a “restricted version” of the proposition: we gave countable double categories over a presheaf category on a countable category, instead of a small double category over a locally presentable category.

### v3

@@ 4.1 Double categories
We will denote $\alpha:l\to r$ for $\alpha$ a square with vertical domain $l$ and codomain $r$. If $\alpha$ is a square of $\mathbb{S}\textbf{q}(\mathcal{C})$ with horizontal domain $u$ and codomain $v$, we will write $\alpha=[u,v]$.
@@ 4.2 Algebraic weak factorisation systems
[…]
@@
    We have vertical compatibilities, which is about composition of vertical morphisms. For any composable vertical morphisms ${A}$${B}$${C}$$\scriptstyle{l}$$\scriptstyle{l^{\prime}}$ in $\mathbb{L}$ and ${X}$${Y}$${Z}$$\scriptstyle{r}$$\scriptstyle{r^{\prime}}$ in $\mathbb{R}$, and any diagram in $\mathcal{C}$ of the following form:
@@
Now we will state without proof Proposition 18 from [5]:
@@
Such lifting structures are called cofibrantly generated. It should be noted that the proof for the above proposition uses a version of Quillen’s small objects argument from [8], where the axiom of choice was used. At the moment, it is not known whether a constructive proof of this result exists.
@@ 4.3 Degenerate-preferring Kan fibrations
${\Lambda^{n}_{m}}$${\Delta^{b}}$${\Delta^{n}}$${\Delta^{b}}$$\scriptstyle{s\circ\iota}$$\scriptstyle{\iota}$$\scriptstyle{id}$$\scriptstyle{s}$ (39)
@@
The left vertical compatibility conditions are trivially satisfied as whenever two vertical arrows can be composed, one of them is an identity. Satisfying the left horizontal compatibility conditions against identity squares is also trivial, thus the only squares for which it is not trivial to satisfy the left horizontal compatibility condition are those as in Diagram 7.
@@
Then $s=s^{\prime}\circ s_{j}$ for some degeneracy map $s_{j}$, hence $z\circ s$ is a degenerate solution for the lifting problem against $\iota$ and therefore $\text{lift}(z\circ s\circ\iota,f\circ s)=z\circ s$ as $\alpha$ is degenerate-preferring. Since $z$ must be the unique lift against the identity arrow, $(\alpha,\text{lift})$ satisfies the left horizontal compatibility condition against the above squares as in Diagram 7 and is a vertical morphism of ${\bf RLP}(\mathbb{D})$.
Conversely, suppose that a degenerate solution $z\circ s_{j}$ to a lifting problem as in Diagram 7 exists, then the square in Diagram 7 factors as the above diagram for $s=s_{j}$. Note that the right square has $z$ as unique lift. By the horizontal compatibility condition, any vertical morphism $(f,\psi)$ of ${\bf RLP}(\mathbb{D})$ therefore assigns $z\circ s_{j}$ as solution to Diagram 7. Therefore any such a vertical morphism is a degenerate-preferring Kan fibration.
@@ 4.4 The double category of horn pushout sequences
In particular, in Diagram 42, the sieve $S_{i}$ contains exactly two more non-degenerate simplices than $S_{i-1}$, namely, one non-degenerate $n$-simplex and one non-degenerate $n-1$-simplex. As a consequence, in Diagram 41, the sieve $S_{k}$ has exactly $2k$ more non-degenerate simplices than $S_{0}$.
For the squares, our first attempt would be the following definition:
@@ 4.5 Functional lifts respect face squares
Note that for any $\iota\in\mathcal{L}$ with $\ell(\iota):\Lambda^{n}_{m}\hookrightarrow\Delta^{n}$, both $\Lambda^{n}_{m}$ and $\Delta^{n}$ are decidable sieves of $\Delta^{n}$ and therefore we have a horn pushout sequence of length $1$ corresponding to $[id,id]:\ell(\iota)\to\ell(\iota)$. If $(f,\psi)$ is a vertical morphism in ${\bf RLP}(\mathbb{S})$, we call its lift against this sequence the lift against $\iota$.
@@
    So let $\delta:\sigma\to\tau$ be a face square and let $\tau$ have length $1$ with chosen $\iota\in\mathcal{L}$. As $\delta$ is a face square, the square $L(\sigma)\to\tau_{1}$ is a pullback square in ${\bf sSet}$ in which all arrows are mono. Therefore, by Remark 4.7, $L(\sigma)$ adds at most 2 faces, thus $\sigma$ has either length 0 or 1. If the length is 00, there is a unique lift against $\sigma$ and we are done.
@@ 4.6 Effective lifts respect degeneracy squares
In this section, we will study what happens if we respect not only face squares, but also degeneracy squares. By Lemma 4.11, we only need to consider degeneracy squares $\delta:\sigma\to\tau$ with $\tau$ of length $1$ as in Diagram 48. Let $s_{j}:\Delta^{b+1}\to\Delta^{b}$ be the underlying simplicial morphism of $\delta$. It is decidable whether the index $j:\Delta^{0}\to\Delta^{b}$ factors through the inclusion $\Delta^{n}\hookrightarrow\Delta^{b}$.
@@
    If $j$ does not factor through $\Delta^{n}$, then the pullback in Diagram 48 restricts to the identity on both $\Lambda^{n}_{m}$ and $\Delta^{n}$ and, therefore, $\Lambda^{n}_{m}$ factors through $s_{j}^{*}(T_{0})$ and $\Delta^{n}$ factors through $s_{j}^{*}(T_{1})$. Therefore, for $(f,\psi)$ a vertical morphism of ${\bf RLP}(\mathbb{S})$, we arrive in the same situation as in Diagram 47, where the vertical domain of $\delta$ has length 1. By Lemma 4.13, both maps $s_{j}^{*}(T_{1})\to X,T_{1}\to X$ are defined by a pushout, and by the universal property of pushouts, both maps $s_{j}^{*}(T_{1})\to X$ are equal, and we respect $\delta$.
@@
      If $j^{\prime}=m$, then $\sigma$ has length 2. The first inclusion $\sigma_{1}$ is the pushout of a horn inclusion $\Lambda^{n}_{m}\hookrightarrow\Delta^{n}$ corresponding to either face $d_{j^{\prime}}$ or $d_{j^{\prime}+1}$ of $\Delta^{n+1}$, while the second inclusion $\sigma_{2}$ is the pushout of the horn inclusion $\Lambda^{n+1}_{m^{*}}\hookrightarrow\Delta^{n+1}$ for $m^{*}$ the face we did not add in $\sigma_{1}$.
@@
      If $j^{\prime}\neq m$, then $\sigma$ has length 3. In this case $\sigma_{1}$ and $\sigma_{2}$ are both the pushout of a horn inclusion $\Lambda^{n}_{m}\hookrightarrow\Delta^{n}$, corresponding to either face $d_{j^{\prime}}$ or $d_{j^{\prime}+1}$ of $\Delta^{n+1}$, while $\sigma_{3}$ is the pushout of the horn inclusion $\Lambda^{n+1}_{m^{*}}\hookrightarrow\Delta^{n+1}$ for $m^{*}=m$ if $m<j^{\prime}$ and $m^{*}=m+1$ if $m>j^{\prime}$.
    For each $\sigma_{i}$ we need a chosen horn inclusion $\iota^{\prime}\in\mathcal{L}$ such that $\sigma_{i}$ is the pushout of $\ell(\iota^{\prime})$. The chosen horn inclusion for the final $\sigma_{i}$ (in the cases above $\sigma_{2}$ and $\sigma_{3}$ respectively) is denoted $\iota^{*}_{\delta}$, for $\iota$ as in Diagram 48. Note that if we respect face squares, the lifts corresponding to the faces $d_{j^{\prime}},d_{j^{\prime}+1}$ of $\Delta^{n+1}$, will be the same as the lift against $\iota$.
@@ 5 Outlook
In Section 2, we have introduced degenerate-preferring and symmetric effective Kan fibrations, both of which can be given the structure of effective Kan fibrations. In [3], the effective Kan fibrations are constructively shown to be closed under pushforwards. This is a first step in a constructive model for HoTT based on simplicial sets. It remains an open question whether these new notions of Kan fibrations are themselves also closed under pushforwards, and can form the basis for another HoTT model.
In Section 3, we have extended the result that all simplicial algebras of an algebraic theory are Kan fibrations iff the algebraic theory is Malcev to our other notions of Kan complexes. As we mentioned in Remark 3.14, there might be an extension of this result to Malcev categories. Another direction for generalization, lies in the Malcev structure in cubical sets with connections. In [22], it has been shown that cubical groups with connections are Kan. It is an open question whether this also holds for cubical Malcev algebras with connections, and whether all algebraic theories whose cubical algebras with connections are Kan are Malcev algebras.
In Section 4, we have presented proofs showing that our new notions fit into lifting algebraic weak factorisation systems. We have used Proposition 4.3, for which it is not yet known whether a constructive proof exists.